Four balls with different masses are dropped from the heights shown.

Air resistance may be ignored.

Which ball has the smallest average speed?


1
Expert's answer
2021-09-16T17:03:12-0400

The ball thrown from the smallest height has the smallest average speed because the average speed for the uniformly accelerated motion is


"u=\frac v2=\frac{\sqrt{2gh}}{2}."